# A Survey on Speech Large Language Models for Understanding

### v5

@@ IX Future Exploration
Building on the challenges discussed in Section VIII., as Speech LLMs continue to evolve, addressing key challenges will be crucial for their progress. Two primary issues that warrant focused research are instruction sensitivity and the degradation of semantic reasoning abilities. Currently, Speech LLMs often struggle to maintain robustness in the face of nuanced or ambiguous instructions. This sensitivity to instruction variability affects their ability to consistently generate accurate and contextually appropriate outputs. Additionally, as Speech LLMs handle increasingly complex tasks, their reasoning capabilities—particularly in understanding subtle, multi-step contexts—tend to degrade. Moreover, Speech LLMs still have a long way to go in achieving true multitask generalizability—where a model can handle tasks it hasn’t been explicitly trained on. At present, models often fail to perform certain tasks unless they have been specifically trained for them. Researchers are expected to focus on improving models’ adaptability to diverse instruction formats, enhancing their reasoning capabilities, and enabling more seamless multitask learning across tasks the model has not been trained on.
Another critical avenue for future work lies in extracting richer and finer acoustic information from speech. While current models can process speech to some extent, their ability to capture the full range of acoustic cues—such as speaker emotion, intonation, and environmental noise—is still limited. Exploring methods to enhance acoustic feature extraction and incorporating these features more effectively into the reasoning process will be key to enabling more nuanced and context-aware speech understanding.
In addition, preference alignment—particularly through Reinforcement Learning from Human Feedback (RLHF)—represents a promising yet underexplored direction in the Speech LLM domain. While RLHF has been widely adopted in the broader LLM community to align model behavior with human intent, its application in Speech LLMs remains relatively limited. As these models are increasingly deployed in real-world, interactive scenarios, it becomes crucial to ensure that their outputs align with user expectations, task-specific requirements, and contextual nuances. RLHF offers a viable mechanism for fine-tuning models based on human preferences and interaction signals, enabling more personalized and context-aware responses [100, 170]. Advancing research in this area could significantly enhance the reliability, controllability, and user satisfaction of Speech LLMs in practical applications.
Together, these areas represent some of the most important directions for future exploration in the development of Speech LLMs. By addressing these challenges, researchers can enhance the utility, adaptability, and overall performance of speech-based models, paving the way for more advanced systems capable of deeper understanding and more effective interaction with users.
@@ X Conclusion
This paper presents the first comprehensive survey of Speech Large Language Models (Speech LLMs) through the lens of speech understanding—a perspective that has been largely overlooked in prior work that has predominantly focused on speech generation or general-purpose multimodal models. To ground this perspective, we propose the first systematic conceptualization of speech understanding, along with a task-oriented taxonomy that spans informational, functional, and format dimensions. Building on this conceptual foundation, we trace the evolution of modeling approaches, highlighting the transition from early modular and cascaded pipelines to End-to-End (E2E) specialized systems and, more recently, to LLM-centric architectures that aim to unify diverse speech understanding tasks. To clarify the current design space, we synthesize recent advances in model architectures, training strategies, and datasets, and further provide an empirical perspective by analyzing how well existing Speech LLMs generalize across a range of speech understanding tasks. Through our own experimental investigation, we identify two pressing challenges in current systems—instruction sensitivity and a degradation in semantic reasoning ability—both of which limit the robustness and flexibility of Speech LLMs in real-world applications. By summarizing the field systematically, identifying key challenges, and proposing actionable directions, we hope this survey can serve as both a conceptual foundation and a practical reference for advancing the development of Speech LLMs toward more general, adaptable, and human-aligned speech understanding systems.

### v6

@@ IX Future Exploration
Building on the challenges discussed in Section VIII, as Speech LLMs continue to evolve, addressing key challenges will be crucial for their progress. Two primary issues that warrant focused research are instruction sensitivity and the degradation of semantic reasoning abilities. Currently, Speech LLMs often struggle to maintain robustness in the face of nuanced or ambiguous instructions. This sensitivity to instruction variability affects their ability to consistently generate accurate and contextually appropriate output. Additionally, as Speech LLMs are applied to increasingly complex tasks, their reasoning capabilities tend to degrade, particularly in understanding subtle, multi-step contexts. Moreover, Speech LLMs still have a long way to go in achieving true multitask generalizability, where a model can handle tasks it hasn’t been explicitly trained on. At present, models often fail to perform certain tasks unless they have been specifically trained for them. Researchers are expected to focus on improving models’ adaptability to diverse instruction formats, enhancing their reasoning capabilities, and enabling more seamless multitask learning across tasks the model has not been trained on.
Another critical avenue for future work lies in extracting richer and finer acoustic information from speech. Although current models can process speech to some extent, their ability to capture the full range of acoustic cues, such as speaker emotion, intonation, and environmental noise, remains limited. Exploring methods to improve the extraction of acoustic information and incorporating these features more effectively into the reasoning process will be key to enabling more nuanced and context-aware speech understanding.
In addition, preference alignment, especially approaches based on reinforcement learning, is a promising direction that has so far received limited attention in Speech LLMs. While it has been widely adopted in the broader LLMs community to align model behavior with human intent, its application in Speech LLMs remains relatively limited. As these models are increasingly deployed in interactive real-world scenarios, it becomes crucial to ensure that their outputs align with user expectations, task-specific requirements, and contextual nuances. RLHF offers a viable mechanism for fine-tuning models based on human preferences and interaction signals, enabling more personalized and context-aware responses [112, 184]. Advancing research in this area could significantly enhance the reliability, controllability, and user satisfaction of Speech LLMs in practical applications.
Together, these areas represent some of the most important directions for future exploration in the development of Speech LLMs. By addressing these challenges, researchers can improve the utility, adaptability, and overall performance of speech-based models, paving the way for more advanced systems capable of deeper understanding and more effective interaction with users.
@@ X Conclusion
This paper presents the first comprehensive survey of Speech LLMs from the perspective of speech understanding. This perspective has been largely overlooked in prior work, which has focused predominantly on speech generation or general-purpose multimodal models. To ground this perspective, we propose the first systematic conceptualization of speech understanding, along with a task-oriented taxonomy that spans informational, functional, and format dimensions. Building on this conceptual foundation, we trace the evolution of modeling approaches, highlighting the transition from early modular and cascaded pipelines to E2E specialized systems, and more recently, to LLM-centric architectures that aim to unify diverse speech understanding tasks. To clarify the current design space, we synthesize recent advances in model architectures, training strategies, and datasets, and further provide an empirical perspective by analyzing how well existing Speech LLMs generalize across a range of speech understanding tasks. Through our own experimental investigation, we identify two pressing challenges: sensitivity to instruction and degradation of semantic reasoning ability of current systems. Both limit the robustness and flexibility of Speech LLMs in real-world applications. By summarizing the field systematically, identifying key challenges, and proposing actionable directions, we hope that this survey can serve as both a conceptual foundation and a practical reference for advancing the development of Speech LLMs toward more general, adaptable, and human-aligned speech understanding systems.